A boxcar contains six complex electronic systems. Two of the six are to be randomly selected for thorough testing and then classified as defective or not defective. (Enter your probabilities as fractions.) (a) If three of the six systems are actually defective, find the probability that at least one of the two systems tested will be defective
Answer #1

P(Both are defective) = (3C3)*(3C0) / (6C3) = 1/20

P(at least one is defective) = P(exactly one defective) + P( both are defective)

P(exactly one defective) =(3C1)*(3C2)) / 6C3= 9/20

P(at least one is defective) = 9/20 + 1/20 = 10/20